Show that the closed sphere with centre (7,3,2)and radius 10 in R³ is contained in the open cube P = {(x, y,z ):|x − 2 |<11,|y− 3|<11, |z − 7 |<11}.


1
Expert's answer
2021-02-23T09:07:08-0500

The equation of the sphere is: "(x-7)^2+(y-3)^2+(z-2)^2=10^2". We point out that the point with coordinates "(17,3,2)" belongs to the sphere. However, the point is not contained in the cube, since "|17-2|=15>11". Thus, the sphere is not contained in in the cube.